You purchased 100 shares of stock for a share price of $16.47. You sold the stock two years later for a share price of $19.33. You also received total dividend payments of $0.72 per share. What was your total return on your investment?
Answer #1
Total return in dollar = Number of shares*(D+(P1-P0))
= 100*(0.72+(19.33-16.47))
= $ 358.00
Where,
D = $       0.72
P0 = $    16.47
P1 = $    19.33
Total return (%) = (D+(P1-P0))/P0
= (0.72+(19.33-16.47))/16.47
= 21.74%
Where,
D = $       0.72
P0 = $    16.47
P1 = $    19.33